# SBP-FDEC: Summation-by-Parts Finite Difference Exterior Calculus

### v2

@@ 4.3 Time Integration
We will consider two different time integration schemes here. One is an explicit three-stage Runge-Kutta scheme [SSPRK] and the other, the discrete gradient average vector field method [discrete_gradient], which in this case coincides with the Crank-Nicolson scheme [crank_nicolson], which is implicit. The purpose of the latter is to exactly preserve the discrete energy in the fully discrete setting. The ansatz for the discrete gradient method is to find an approximation for $\overline{\nabla}_{U}H(\underline{U}_{n},\underline{U}_{n+1})$, that depends on the current and the next time step, for $\nabla H(\underline{U})$. This approximation has to fulfill
@@
If we then define the time integration scheme implicitly by
@@
which conserves the Hamiltonian [discrete_gradient] since
@@
Since $\nabla_{U}H$ is linear, this choice leads to the Crank-Nicolson time integration method [crank_nicolson]:
@@
where we used (102). For the Crank-Nicolson method we have
@@ 5 Numerical Verification
We look at the test case from [sonnendruecker_maxwell_test_case] with domain $\Omega=[-1,1]^{2}$, end time $T=1$, and the initial condition
@@
$E^{x}(x,y)$ $=-\frac{1}{\sqrt{2}}\text{cos}(\pi x+\pi)\text{sin}(\pi y+\pi)\text{sin}(\sqrt{2}\pi t),$ (132)
$E^{y}(x,y)$ $=\frac{1}{\sqrt{2}}\text{sin}(\pi x+\pi)\text{cos}(\pi y+\pi)\text{sin}(\sqrt{2}\pi t),$ (133)
$B^{z}(x,y)$ $=\text{cos}(\pi x+\pi)\text{cos}(\pi y+\pi)\text{cos}(\sqrt{2}\pi t).$ (134)
For our tests we use a Julia code, using scripts from [fernandez2014review]. To solve the linear system for the Crank-Nicolson method, we use the GMRES solver [gmres] implemented in the Julia package Krylov.jl [montoison-orban-2023]. For the repository, see the data availability section. The tests are run with Julia 1.11.7 on Ubuntu 24.04.3 LTS with a AMD Ryzen 9 9900X processor.
@@ 5.1 Convergence Tests
We run two different kinds of convergence tests, where we use the discrete $L^{2}$ norm induced by our global mass matrices. For the first kind of convergence test we keep the number of points per element constant but varying the amount of elements. For time integration we use a three-stage strong stability-preserving Runge-Kutta scheme (see [SSPRK], equation (2.18)) with a constant time step of $\Delta t=2\cdot 10^{-5}$. We run this test for operators of boundary order $2$ and $3$ by Strand [strand_sbp] several times, varying the amount of points per element between test sets, to both include operators containing only the boundary stencils and operators mostly containing inner stencils.
@@
For both polynomial degrees we observe an even-odd behavior in figures 6-11, where we obtain an alternating order of $p$ and $p+1$ for the electric field, depending on the number of points per element. This behavior is documented for central schemes, see for instance [even_odd_convergence]. For the magnetic field the order is more erratic, nonetheless do we obtain an average EOC close to $p+1$. The second kind of test is again run using operators from [strand_sbp] of polynomial degrees $2$ and $3$. The time integration this time is now performed using the same SSPRK scheme from the first tests, but also with the energy-conservative Crank-Nicolson scheme. We again use a constant time step of $\Delta t=2\cdot 10^{-5}$. For the spatial discretization we just use one element, but a varying amount of nodes per element. In figures 12-15 we see higher polynomial orders than in the first test, which is likely due to the fact that the inner stencils, which are of degree $2p$ become more influential and increase the EOC.
@@ 5.2 Divergence Tests
A key property of the proposed discretization is the exact preservation of the divergence condition (55). We test this using both a coarse discretization using $2$ elements with $12$ points and a finer discretization using $5$ elements with $20$ points in each direction. For the coarse discretization we also simulate up to the end time $T=10000$. We again test the divergence preservation with the operators of degree $2$ and $3$ and both with the SSPRK and the Crank-Nicolson scheme. The time step is defined by
@@
for $\text{CFL}=1$. Here $\omega_{min}$ is the minimal quadrature weight of the global mass matrix $\hat{M}_{2D}$. The figures 16, 17 and 18 show the maximum absolute value of the divergence of the electric field at the grid points. We can observe that the solution remains divergence free up to machine precision during the entire simulation, regardless of operator, time integration scheme, or grid size. The rounding errors for the SSPRK scheme do not increase anymore after a certain time, which is due to the fact that the energy has almost completely dissipated away. This confirms the previously proven fully discrete divergence preservation.
@@ 5.3 Energy Conservation Tests
We repeat the tests performed for divergence preservation to test the conservation of energy over time. We use the same operators and the same time integration schemes with the grid sizes. In figures 19, 20 and 21 we plot the difference between the discrete energy at $t=0$ and the energy at the given point in time. We can observe that the SSPRK is energy dissipative and does not conserve the energy exactly. This effect lessens for the finer grid size, but is still present. On the other hand the Crank Nicolson time integration method conserves the energy exactly as proven previously.
Number of Elements $E^{x}/E^{y}$ Error $E^{x}/E^{y}$ EOC $B^{z}$ Error $B^{z}$ EOC
1 1.487802e11.487802e-1 - 2.128334e12.128334e-1 -
2 1.457635e21.457635e-2 $3.35$ 3.770852e23.770852e-2 $2.50$
4 3.215945e33.215945e-3 $2.18$ 2.840422e32.840422e-3 $3.73$
8 4.087012e44.087012e-4 $2.98$ 2.402008e42.402008e-4 $3.56$
16 4.357807e54.357807e-5 $3.23$ 4.276273e54.276273e-5 $2.49$
32 5.590312e65.590312e-6 $2.96$ 8.925102e68.925102e-6 $2.26$
Figure 6: Discrete $L^{2}$-Errors and EOC for degree $p=2$ with $8$ points per direction per element
Number of Elements $E^{x}/E^{y}$ Error $E^{x}/E^{y}$ EOC $B^{z}$ Error $B^{z}$ EOC
1 9.964855e29.964855e-2 - 1.260615e11.260615e-1 -
2 1.130372e21.130372e-2 $3.14$ 5.891993e35.891993e-3 $4.42$
4 2.212383e32.212383e-3 $2.35$ 1.149555e31.149555e-3 $2.36$
8 5.62475e45.62475e-4 $1.98$ 1.331211e41.331211e-4 $3.11$
16 1.407649e41.407649e-4 22 5.03316965.033169-6 $4.73$
32 3.518716e53.518716e-5 22 1.586997e61.586997e-6 $1.67$
Figure 7: Discrete $L^{2}$-Errors and EOC for degree $p=2$ with $9$ points per direction per element
Number of Elements $E^{x}/E^{y}$ Error $E^{x}/E^{y}$ EOC $B^{z}$ Error $B^{z}$ EOC
1 4.418904e24.418904e-2 - 1.054903e11.054903e-1 -
2 3.740646e33.740646e-3 $3.56$ 1.440109e21.440109e-2 $2.87$
4 1.34654e31.34654e-3 $1.47$ 1.170316e31.170316e-3 $3.62$
8 1.722958e41.722958e-4 $2.97$ 2.350005e42.350005e-4 $2.32$
16 2.678038e52.678038e-5 $2.69$ 1.035082e51.035082e-5 $4.5$
32 2.699576e62.699576e-6 $3.31$ 2.37598962.375989-6 $2.12$
Figure 8: Discrete $L^{2}$-Errors and EOC for degree $p=2$ with $10$ points per direction per element
@@
32 2.215563e72.215563e-7 $3.02$ 2.44092782.440927-8 $4.51$
Figure 9: Discrete $L^{2}$-Errors and EOC for degree $p=3$ with $12$ points per direction per element
@@
Figure 10: Discrete $L^{2}$-Errors and EOC for degree $p=3$ with $13$ points per direction per element
Number of Elements $E^{x}/E^{y}$ Error $E^{x}/E^{y}$ EOC $B^{z}$ Error $B^{z}$ EOC
1 8.506095e38.506095e-3 - 2.235792e22.235792e-2 -
2 9.643637e49.643637e-4 $3.14$ 3.285374e33.285374e-3 $2.77$
4 1.017081e41.017081e-4 $3.25$ 4.232642e54.232642e-5 $6.28$
8 7.697999e67.697999e-6 $3.72$ 4.7801012e64.7801012e-6 $3.15$
16 9.530514e79.530514e-7 $3.01$ 1.958355e71.958355e-7 $4.61$
32 1.176027e71.176027e-7 $3.02$ 1.007924e81.007924e-8 $4.28$
Figure 11: Discrete $L^{2}$-Errors and EOC for degree $p=3$ with $14$ points per direction per element
Points per Element $E^{x}/E^{y}$ Error $E^{x}/E^{y}$ EOC $B^{z}$ Error $B^{z}$ EOC
8 1.487802e11.487802e-1 - 2.128334e12.128334e-1 -
16 9.078998e39.078998e-3 $4.03$ 1.586406e21.586406e-2 $3.75$
32 4.807838e44.807838e-4 $4.24$ 8.52153e48.52153e-4 $4.22$
64 3.536512e53.536512e-5 $3.76$ 4.897816e54.897816e-5 $4.12$
128 2.885322e62.885322e-6 $3.62$ 2.921983e62.921983e-6 $4.07$
256 2.459519e72.459519e-7 $3.55$ 1.785489e71.785489e-7 $4.03$
Figure 12: Discrete $L^{2}$-Errors and EOC for degree $p=2$ with SSPRK time integration with varying points per element
@@
Figure 13: Discrete $L^{2}$-Errors and EOC for degree $p=3$ with SSPRK time integration with varying points per element
@@
8 1.487802e11.487802e-1 - 2.128334e12.128334e-1 -
16 9.078998e39.078998e-3 $4.03$ 1.586407e21.586407e-2 $3.75$
32 4.807839e44.807839e-4 $4.24$ 8.521539e48.521539e-4 $4.22$
64 3.536516e53.536516e-5 $3.76$ 4.897902e54.897902e-5 $4.12$
128 2.885341e62.885341e-6 $3.62$ 2.92283262.922832-6 $4.07$
256 2.459736e72.459736e-7 $3.55$ 1.794117e71.794117e-7 $4.03$
Figure 14: Discrete $L^{2}$-Errors and EOC for degree $p=2$ with Crank-Nicolson time integration with varying points per element
@@
12 1.566639e21.566639e-2 - 2.98044e22.98044e-2 -
24 1.1644431.16444-3 $3.75$ 3.248053e33.248053e-3 $3.2$
48 9.278422e59.278422e-5 $3.65$ 1.437682e41.437682e-4 $4.5$
96 5.909218e65.909218e-6 $3.97$ 6.813825e66.813825e-6 $4.4$
192 3.838071e73.838071e-7 $3.94$ 2.91974572.919745-7 $4.54$
384 2.34832182.348321-8 $4.03$ 1.33955781.339557-8 $4.45$
Figure 15: Discrete $L^{2}$-Errors and EOC for degree $p=3$ with Crank-Nicolson time integration with varying points per element
Figure 16: Divergence errors for operators of degree $p=2,3$ with both the SSPRK and Crank-Nicolson time integration, 22 elements with $12$ points per dimension
Figure 17: Divergence errors for operators of degree $p=2,3$ with both the SSPRK and Crank-Nicolson time integration, 55 elements with $20$ points per dimension
Figure 18: Divergence errors for operators of degree $p=2,3$ with both the SSPRK and Crank-Nicolson time integration, 22 elements with $12$ points per dimension, end time $T=10000$
Figure 19: Energy errors for operators of degree $p=2,3$ with both the SSPRK and Crank-Nicolson time integration, 22 elements with $12$ points per dimension
Figure 20: Energy errors for operators of degree $p=2,3$ with both the SSPRK and Crank-Nicolson time integration, 55 elements with $20$ points per dimension
Figure 21: Energy errors for operators of degree $p=2,3$ with both the SSPRK and Crank-Nicolson integration, 22 elements with $12$ points per dimension, end time $T=10000$

### v3

@@ 4.3 Time Integration
We will consider two different time integration schemes here. One is an explicit three-stage Runge-Kutta scheme [43] and the other, the discrete gradient average vector field method [31], which in this case coincides with the Crank-Nicolson scheme [23], which is implicit. The purpose of the latter is to exactly preserve the discrete energy in the fully discrete setting. The ansatz for the discrete gradient method is to find an approximation $\overline{\nabla}_{U}H(\underline{U}_{n},\underline{U}_{n+1})$, that depends on the current and the next time step, for $\nabla H(\underline{U})$. We require this approximation to fulfill
@@
because definining the time integration scheme implicitly by
@@
conserves the Hamiltonian [31] since
@@
Since $\nabla_{U}H$ is linear, this choice leads to the Crank-Nicolson time integration method [23]:
@@
where we used (115) and
$\begin{pmatrix}\underline{A}\\ \underline{B}\end{pmatrix}\cdot\begin{pmatrix}\underline{v}\\ \underline{w}\end{pmatrix}\vcentcolon=\begin{pmatrix}\underline{A}\underline{v}\\ \underline{B}\underline{w}\end{pmatrix}$, (152)
for matrices $\underline{A},\underline{B}$ and vectors $\underline{v},\underline{w}$ with compatible dimensions. For the Crank-Nicolson method we have
@@ 5 Numerical Verification
We look at two test cases. The first is from [39] with domain $\Omega=[-1,1]^{2}$, end time $T=1$, periodic boundary conditions, and the initial condition
@@
$E^{x}(x,y,t)$ $=-\frac{1}{\sqrt{2}}\text{cos}(\pi x+\pi)\text{sin}(\pi y+\pi)\text{sin}(\sqrt{2}\pi t),$ (159)
$E^{y}(x,y,t)$ $=\frac{1}{\sqrt{2}}\text{sin}(\pi x+\pi)\text{cos}(\pi y+\pi)\text{sin}(\sqrt{2}\pi t),$ (160)
$B^{z}(x,y,t)$ $=\text{cos}(\pi x+\pi)\text{cos}(\pi y+\pi)\text{cos}(\sqrt{2}\pi t).$ (161)
The second test case has domain $\Omega=[0,1]^{2}$, end time $T=1$, initial condition
$E^{x}(x,y)$ $=0,$ (162)
$E^{y}(x,y)$ $=0,$ (163)
$B^{z}(x,y)$ $=\sqrt{2}\text{cos}(\pi x)\text{sin}(\pi y),$ (164)
with essential homogeneous boundary condition for $y=0$ and $y=1$, and natural boundary conditions for $x=0$ and $x=1$, resulting in:
$E^{y}(x,y,t)$ $=0\text{ on }\partial\Omega\times[0,1],$ (165)
$B^{z}(0,y,t)$ $\equiv 0,$ (166)
$B^{z}(1,y,t)$ $\equiv 0.$ (167)
This problem has the analytical solution
$E^{x}(x,y,t)$ $=\text{cos}(\pi x)\text{cos}(\pi y)\text{sin}(\sqrt{2}\pi t),$ (168)
$E^{y}(x,y,t)$ $=\text{sin}(\pi x)\text{sin}(\pi y)\text{sin}(\sqrt{2}\pi t),$ (169)
$B^{z}(x,y,t)$ $=\sqrt{2}\text{cos}(\pi x)\text{sin}(\pi y)\text{cos}(\sqrt{2}\pi t).$ (170)
For our tests we use a Julia code, using scripts from [11]. To solve the linear system for the Crank-Nicolson method, we use the GMRES solver [41] implemented in the Julia package Krylov.jl [33]. For the repository, see the data availability section. The tests are run with Julia 1.11.7 on Ubuntu 24.04.4 LTS with a AMD Ryzen 9 9900X processor.
@@ 5.1 Convergence Tests
We run two different kinds of convergence tests, where we use the discrete $L^{2}$ norm induced by our global mass matrices. This means we compute the error between the nodal values of the numerical approximation $\underline{E}^{x}_{\text{num}}$, $\underline{E}^{y}_{\text{num}}$, $\underline{B}^{z}_{\text{num}}$ and the exact nodal values at time $T=1$, $\underline{E}^{x}$, $\underline{E}^{y}$, $\underline{B}^{z}$ as follows:
$\|\underline{E}^{x}_{\text{num}}-\underline{E}^{x}\|_{L^{2},E_{x},num}$ $\vcentcolon=\sqrt{\left(\underline{E}^{x}_{\text{num}}-\underline{E}^{x}\right)^{T}\underline{M}^{2D}_{x}\left(\underline{E}^{x}_{\text{num}}-\underline{E}^{x}\right)},$ (171)
$\|\underline{E}^{y}_{\text{num}}-\underline{E}^{y}\|_{L^{2},E_{y},num}$ $\vcentcolon=\sqrt{\left(\underline{E}^{y}_{\text{num}}-\underline{E}^{y}\right)^{T}\underline{M}^{2D}_{y}\left(\underline{E}^{y}_{\text{num}}-\underline{E}^{y}\right)},$ (172)
$\|\underline{B}^{z}_{\text{num}}-\underline{B}^{z}\|_{L^{2},B^{z},num}$ $\vcentcolon=\sqrt{\left(\underline{B}^{z}_{\text{num}}-\underline{B}^{z}\right)^{T}\underline{\hat{M}}^{2D}\left(\underline{B}^{z}_{\text{num}}-\underline{B}^{z}\right)}.$ (173)
@@
For the first kind of convergence test we keep the number of points per element constant but vary the amount of elements. For time integration we use a three-stage strong stability-preserving Runge-Kutta scheme (see [43], equation (2.18)) with a constant time step of $\Delta t=2\cdot 10^{-5}$. We run this test for the operator of boundary order $3$ by Strand [44] for 12 and 13 points per element in the periodic case and for 12 points again for the non-periodic case to investigate the behavior of operators dominated by boundary stencils. Appendix B contains further results the periodic case for 14 points per element and similar results for the operator of boundary order $2$ by Strand [44]. Because of the symmetry of the example, the errors for $E^{x}$ and $E^{y}$ are identical up to machine precision and are therefore only listed once in the periodic case.
We observe an even-odd behavior in Tables 1 and 2 and further results in appendix B where we obtain an alternating order of $p$ and $p+1$ for the electric field, depending on the number of points per element. This behavior is reminiscent of the even-odd behavior previously reported for central (non-dissipative) continuous and discontinuous Galerkin discretization, see for instance [19] and [1]. However, the analogy should be interpreted carefully. In the Galerkin methods, the discretization is parameterized by the polynomial degree, whereas the present formulation is defined directly through nodal SBP finite-difference operators and therefore has no underlying polynomial-degree parameter. Consequently, the analogous behavior appears naturally in the number of nodal degrees of freedom per element.
As detailed above, present method employs a continuous-Galerkin-type superposition at element interfaces rather than dissipative SAT penalties or upwind numerical fluxes. The resulting discretization is therefore essentially non-dissipative, similarly to central continuous and discontinuous Galerkin formulations where parity-dependent convergence behavior has previously been observed [19, 1]. While we are not aware of previous studies reporting the exact same behavior for central SBP finite-difference discretizations, related even-odd effects have been reported for SBP finite-difference operators in connection with the treatment of high-frequency numerical modes [30, 12]. This suggests that the observed behavior may be linked to the discrete high-frequency mode structure of the nodal SBP operators and the absence of interface dissipation, rather than to the formal truncation order alone.
The magnetic field EOC varies without clearly converging towards a set number, though the error of the magnetic field tends to improve at least as good if not faster then the electric field error. The non-periodic case for $12$ points in table 6 shows the same behavior as the periodic case.
@@
To further investigate the behavior of the magnetic field convergence, we ran additional simulations for the non-periodic case with the $p=3$ operator with a varying amount of elements. We ran the problem to end time $T=0.1$ with element numbers per direction from $21$ to $120$ elements. Figure 9 shows a log-log plot of element width $h$ against the discrete L2-error of the magnetic field. We see that the convergence of the magnetic field is bounded from above by a polynomial of degree $4$, but has a highly non-monotonous convergence behavior, exhibiting a periodic pattern (when seen as a function in $h^{-1}$). Further analysis of this behavior is beyond the scope of this paper, but we see that at least in this case we have a convergence behavior that is still $\mathcal{O}(h^{4})$. The second kind of test is again run using operators from [44] of polynomial degree $3$ for both the periodic and the non-periodic case. We again use the SSPRK time integration scheme, but for the periodic case, further results using the Crank-Nicolson time integration scheme and also with the $p=2$ operator can be found in appendix B and exhibit similar behavior. We again use a constant time step of $\Delta t=2\cdot 10^{-5}$. For the spatial discretization we just use one element, but a varying number of nodes per element. In tables 3 and 7 we see higher polynomial orders than in the first test, which is likely due to the fact that the inner stencils, which are of degree $2p$ become more influential and increase the EOC.
We also ran this convergence test for the periodic test case with a non-SBP operator using the mass matrix of the $p=2$ and the derivative matrix of the $p=3$ operator, which together do not have the SBP property. Since strong and weak form of the scheme are no longer equivalent for this operator, we run a convergence test for both the strong and the weak form of the scheme. Table 4 shows the convergence rates for the strong from, exhibiting similar convergence rates to the $p=3$ SBP operator test case, while the convergence rates of the weak form variant in table 5 are significantly lower.
@@ 5.2 Divergence Tests
A key property of the proposed discretization is the exact preservation of the divergence condition (66). We test this using both a discretization using $5$ elements with $20$ points in each direction for both the periodic and the non-periodic test case. We test divergence preservation with the operators of degree $2$ and $3$ and both with the SSPRK and the Crank-Nicolson scheme. The time step is defined by
@@
for $\text{CFL}=1$. Here $\omega_{min}$ is the minimal quadrature weight of the global mass matrix $\hat{M}_{2D}$. Because we assume unit light speed, the maximum absolute value of the wave speeds is 11 and therefore does not appear in (174). The figures 10 and 13 show the maximum absolute value of the divergence of the electric field at the grid points. We can observe that the solution remains divergence free up to machine precision during the entire simulation, regardless of operator, time integration scheme, or grid size. This confirms the previously proven fully discrete divergence preservation.
@@ 5.3 Energy Conservation Tests
We repeat the tests performed for divergence preservation to test the conservation of energy over time. We use the same operators and the same time integration schemes with the grid sizes. In figures 11 and 14 we plot the difference between the discrete energy at $T=0$ and the energy at the given point in time. We can observe that the SSPRK is energy dissipative and does not conserve the energy exactly. This effect lessens for the finer grid size, but is still present. On the other hand the Crank Nicolson time integration method conserves the energy exactly as proven previously. We also run an additional test for the periodic case for energy conservation of the non-SBP operator with mass matrix from the $p=2$ operator and the derivative matrix from the $p=3$ operator for both the strong and the weak form of the scheme using $2$ elements and $12$ points per element in each dimension. To make the influence of the time integration scheme negligible, we use the a time step of $\Delta t=2\cdot 10^{-5}$ with SSPRK time integration. Figure 12 shows that the weak form preserves energy, but the strong form does not do so. The maximum absolute error for the weak form due to time integration is about $1.5\cdot 10^{-13}$. This suggests that for non-SBP operators the weak form is energy conservative, but has a worse convergence behavior, while the strong form has expected convergence rates but is not energy-conservative.
@@
32 2.215563e72.215563e-7 $3.02$ 2.440927e82.440927e-8 $4.51$
Table 1: Discrete $L^{2}$-Errors and EOC for the periodic problem for degree $p=3$ with $12$ points per direction per element
@@
Table 2: Discrete $L^{2}$-Errors and EOC for the periodic problem for degree $p=3$ with $13$ points per direction per element
@@
Table 3: Discrete $L^{2}$-Errors and EOC for the periodic problem for degree $p=3$ with SSPRK time integration with varying points per element with one element in total
@@
12 1.541206e21.541206e-2 - 3.321995e23.321995e-2 -
24 1.284108e31.284108e-3 $3.59$ 3.362758e33.362758e-3 $3.3$
48 9.654045e59.654045e-5 $3.73$ 1.479463e41.479463e-4 $4.51$
96 6.080192e66.080192e-6 $3.99$ 7.090472e67.090472e-6 $4.38$
192 3.978717e73.978717e-7 $3.93$ 3.070251e73.070251e-7 $4.53$
384 2.391862e82.391862e-8 $4.06$ 1.370273e81.370273e-8 $4.49$
Table 4: Discrete $L^{2}$-Errors and EOC for the periodic problem for the non-SBP operator in the strong form with SSPRK time integration with varying points per dimension per element with one element in total
@@
12 1.840316e11.840316e-1 - 1.938346e11.938346e-1 -
24 9.051131e29.051131e-2 $1.02$ 9.118747e29.118747e-2 $1.09$
48 2.684118e22.684118e-2 $1.75$ 1.223276e21.223276e-2 $2.9$
96 9.0215683e39.0215683e-3 $1.57$ 2.224402e32.224402e-3 $2.46$
192 3.3507205e33.3507205e-3 $1.43$ 3.815636e43.815636e-4 $2.54$
384 1.1464041e31.1464041e-3 $1.55$ 4.608534e54.608534e-5 $3.05$
Table 5: Discrete $L^{2}$-Errors and EOC for the periodic problem for the non-SBP operator in the weak form with SSPRK time integration with varying points per element with one element in total
Elements $E^{x}$ Error $E^{x}$ EOC $E^{y}$ Error $E^{y}$ EOC $B^{z}$ Error $B^{z}$ EOC
1 1.6079e31.6079e-3 - 1.8418e31.8418e-3 - 2.8864e32.8864e-3 -
2 1.679e41.679e-4 $3.26$ 1.679e41.679e-4 $3.46$ 4.7262e54.7262e-5 $5.93$
4 1.1008e51.1008e-5 $3.93$ 1.1008e51.1008e-5 $3.93$ 3.7225e63.7225e-6 $3.67$
8 1.2695e61.2695e-6 $3.12$ 1.2695e61.2695e-6 $3.12$ 3.942e73.942e-7 $3.24$
16 1.5666e71.5666e-7 $3.02$ 1.5666e71.5666e-7 $3.02$ 1.726e81.726e-8 $4.51$
32 1.9483e81.9483e-8 $3.01$ 1.9483e81.9483e-8 $3.01$ 7.5131e107.5131e-10 $4.52$
Table 6: Discrete $L^{2}$-Errors and EOC for the non-periodic problem for degree $p=3$ with $12$ points per direction per element
Points $E^{x}$ Error $E^{x}$ EOC $E^{y}$ Error $E^{y}$ EOC $B^{z}$ Error $B^{z}$ EOC
12 1.6079e31.6079e-3 - 1.8418e31.8418e-3 - 2.8864e32.8864e-3 -
24 7.7133e57.7133e-5 $4.38$ 5.4729e55.4729e-5 $5.07$ 1.6725e41.6725e-4 $4.11$
48 5.0039e65.0039e-6 $3.95$ 3.7549e63.7549e-6 $3.87$ 7.6925e67.6925e-6 $4.44$
96 3.136e73.136e-7 $4.0$ 2.8876e72.8876e-7 $3.7$ 3.5209e73.5209e-7 $4.45$
192 1.9237e81.9237e-8 $4.03$ 2.0145e82.0145e-8 $3.84$ 1.5232e81.5232e-8 $4.53$
384 1.1712e91.1712e-9 $4.04$ 1.259891.2598-9 $4.0$ 7.5632e107.5632e-10 $4.33$
Table 7: Discrete $L^{2}$-Errors and EOC for the non-periodic problem for degree $p=3$ with SSPRK time integration with varying points per element per dimension with one element in total
Figure 9: Log-log plot of the magnetic field error against the element size for the non-periodic problem for degree $p=3$ with SSPRK time integration with element number between 2121 and $120$ elements
Figure 10: Divergence errors for the periodic problem for operators of degree $p=2,3$ with both the SSPRK and Crank-Nicolson time integration, 55 elements with $20$ points per dimension
Figure 11: Energy errors for the periodic problem for operators of degree $p=2,3$ with both the SSPRK and Crank-Nicolson time integration, 55 elements with $20$ points per dimension
Figure 12: Energy errors for the periodic problem for the non-SBP operator in strong and weak form with SSPRK time integration, 22 elements with $12$ points per dimension. The maximum absolute energy error for the weak form due to time integration is  $1.5\cdot 10^{-13}$.
Figure 13: Divergence errors for the non-periodic problem for operators of degree $p=2,3$ with both the SSPRK and Crank-Nicolson time integration, 55 elements with $20$ points per dimension
Figure 14: Energy errors for the non-periodic problem for operators of degree $p=2,3$ with both the SSPRK and Crank-Nicolson time integration, 55 elements with $20$ points per dimension